OpenStudy (anonymous):

what is the solution to the inequality 4 times a is less than or equal to twelve?

OpenStudy (anonymous):

To isolate a, you must "undo" the 4 times. So divide both sides of the inequality by 4 to get your answer.

OpenStudy (anonymous):

I tried that and somehow I got15

OpenStudy (jdoe0001):

hmmm what would you get if you were to solve the EQUALity of 4a = 12 ?

OpenStudy (anonymous):

You have\[4a \le 12\]Divide both sides by 4, giving\[\frac{ 4a }{ 4 }\le \frac{ 12 }{ 4 }\]Can you finish it off?

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

thx

OpenStudy (anonymous):

Great!